OpenStudy (math2400):

determine the sum of this series...i thought it was divergent but I guess it diverges cuz it says that it's not divergent

OpenStudy (math2400):

\[\sum_{n=1}^{\infty} \frac{ (-3)^{n-1} }{ 5^n}\]

OpenStudy (math2400):

@jim_thompson5910 would u be able to help me on this please?

ganeshie8 (ganeshie8):

As a start, write the fraction in single exponent

ganeshie8 (ganeshie8):

\[\sum_{n=1}^{\infty} \frac{ (-3)^{n-1} }{ 5^n} = \sum_{n=1}^{\infty} \frac{ (-3)^{n} }{ (-3)5^n} = -\frac{1}{3}\sum_{n=1}^{\infty} \left(\frac{ -3 }{ 5}\right)^n\]

ganeshie8 (ganeshie8):

next recall geometric series

OpenStudy (math2400):

a/(1 - r) ??

ganeshie8 (ganeshie8):

Yep

OpenStudy (math2400):

what would be r? this is where I'm confused

ganeshie8 (ganeshie8):

r = the stuff under the exponent = -3/5

OpenStudy (math2400):

and for a i just plug in 1 into the equation?

OpenStudy (math2400):

and get that value?

ganeshie8 (ganeshie8):

there is no equation

ganeshie8 (ganeshie8):

for "a", just plugin n=1 into the general term of series

OpenStudy (math2400):

that's what i mean haha :)

ganeshie8 (ganeshie8):

i kno, just making sure you see the difference between "equation" and "expression"

OpenStudy (math2400):

okay i plugged everything in but it's zero??

ganeshie8 (ganeshie8):

what do you get for first term, a ?

OpenStudy (math2400):

(-3)^(1-1)/5^1=0

ganeshie8 (ganeshie8):

Ah no, below is our slightly massaged series : \[-\frac{1}{3}\color{blue}{\sum_{n=1}^{\infty} \left(\frac{ -3 }{ 5}\right)^n}\] plugin \(n=1\), you get \(a_1 = -\frac{3}{5} \) \(r = -\frac{3}{5}\) so the infinite sum is \[\dfrac{-3/5}{1-(-3/5)}\] simplify

OpenStudy (math2400):

-0.375

OpenStudy (math2400):

that's still not right tho haha

ganeshie8 (ganeshie8):

looks good, don't forget that -1/3 in front

OpenStudy (math2400):

got it! thanks :)

ganeshie8 (ganeshie8):

np